OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point 4x-12y=2; (10,-1)

OpenStudy (anonymous):

ok, so first my advice would be to put your line equation into standard form. y = mx+b

OpenStudy (anonymous):

can you help me with that

OpenStudy (anonymous):

yeah, 4x-12y = 2 -12y=2-4x y = (2/-12)-(4/-12)x

OpenStudy (anonymous):

okay thank you and what is next?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yeah one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok, so we now know that the equation of the first line is y = (-1/3)x -6. to find the perpendicular line you have to find the negative reciprocal of the slope -1/3 to +3 however, since the line has to go through a specific point, the y intercept (-6) changes. therefore, you now have y = 3x+b to find b, you plug in the values of (10,-1) for (x,y). therefore, 10 = 3(-1) + b 10+3 =b b = 13 therefore, y = 3x +13 is the equation of the line perpendicular to your first line which also goes through (10,-1)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw